# Cost-aware LLM-based Online Dataset Annotation

### v1

@@ 1 Introduction
The rapid proliferation of data across domains has created an urgent need for accurate, large-scale annotation pipelines. While human experts and crowd workers have been the gold standard for dataset labeling, manual annotation is notoriously slow, expensive, and prone to inter-annotator inconsistency Petrović et al. (2020). As machine learning models become increasingly sophisticated, their demand for high-quality, richly labeled datasets only intensifies, exacerbating this bottleneck.
Recent advances in large language models (LLMs) offer a promising remedy: by leveraging transformer-based architectures such as GPT, it is now possible to automate much of the labeling workload. LLMs excel at natural-language understanding, reasoning, and contextual inference, enabling rapid generation of annotations with minimal human effort Naveed et al. (2023).
However, relying on a single LLM introduces issues of biases inherited from its training data and stochastic variability across repeated queries, undermining reliability and reproducibility Errica et al. (2024); Li et al. (2024a). A common strategy to bolster label quality is ensembling: querying multiple LLMs, or multiple samples from the same model, and aggregating their outputs via majority voting. This reduces hallucinations and offsets the bias of individual models but substantially increases cost, as each additional model increases latency and compute expenditure Yang et al. (2023). In practice, querying every available LLM for each instance is often wasteful and unnecessary.
In this paper, we address this trade-off by adaptively selecting a subset of LLMs for majority voting on each input, achieving comparable accuracy to full-ensemble voting while dramatically cutting cost. Unlike prior work on LLM weight optimization or query routing—which presumes access to ground-truth labels or a pre-trained routing model Chen et al. (2024); Nguyen et al. (2024); Ding et al. (2024)—our method operates online, without any held-out training set or ground truth.
@@
    Cost-aware Majority Voting (CaMVo). We propose CaMVo, an algorithm that combines a LinUCB-style contextual bandit with a Bayesian Beta-mixture confidence estimator. For each candidate LLM, CaMVo computes a lower confidence bound on its correctness probability given the input’s embedding, then selects the smallest-cost subset whose aggregated confidence exceeds a user-specified threshold $\delta$.
@@ 1.1 Related Work
Aggregating outputs from multiple LLMs (or repeated queries to a single LLM) via majority voting has become a popular strategy to boost annotation reliability. Chen et al. (2024) analyze the effect of repeated queries to a single model and observe a non-monotonic accuracy curve: performance improves initially but degrades beyond an optimal number of calls due to task heterogeneity. While additional LM calls enhance accuracy on easier queries, they may introduce noise or inconsistency that degrades performance on more challenging ones. To address this, the authors propose a scaling model that predicts the optimal number of LM calls required to maximize aggregate performance. Trad and Chehab (2024) compare re-querying and multi-model ensembles, showing that ensemble strategies are most effective when individual models or prompts exhibit comparable performance levels, and ensemble gains may diminish when individual model accuracies diverge.
Yang et al. (2023) propose a weighted majority-voting ensemble for medical QA, combining dynamic weight adjustment with clustering-based model selection. However, their approach relies on offline training data and focuses solely on improving accuracy, without accounting for the cost of querying. In contrast, our approach selects a cost-effective subset of heterogeneous LLMs online, without any pre-training.
@@
Query routing addresses the problem of selecting a single LLM per query to optimize cost or latency under performance constraints. Nguyen et al. (2024) cast LLM selection as a contextual bandit problem, training offline on labeled data to learn a routing policy that maps query embeddings to a single optimal LLM under a total budget constraint of $b$. Ding et al. (2024) train a router to distinguish “easy” versus “hard” queries, sending easy tasks to local LLMs and hard ones to cloud APIs. Unlike these methods, our method operates in a fully online setting without access to labeled training data, updating the model dynamically during the labeling process. Moreover, rather than routing to a single LLM, we select a cost-efficient subset for majority voting.
@@
Estimating model confidence can guide automatic label selection. Kadavath et al. (2022) explore how a language model’s own uncertainty estimates can serve as predictors of answer correctness, and find that the cumulative log-probability the model assigns to its generated token sequence correlates strongly with factual accuracy across diverse benchmarks. Li et al. (2024a) generate code multiple times and use output similarity as a proxy for confidence, choosing the most consistent result. While these works focus on self-consistency of a single model, we estimate a probabilistic lower bound on each LLM’s correctness via a Bayesian Beta-mixture model, incorporating both past performance and context.
@@
Beyond simple voting, weighted schemes assign each annotator or model a reliability score such as the accuracy of the annotator, or the label confidence reported by the annotator. One notable approach is the GLAD model by Whitehill et al. (2009), which formulates weighted majority voting as a probabilistic inference problem over annotator expertise and task difficulty. GLAD jointly estimates per-annotator reliability and per-item ambiguity via a generative model, using an EM algorithm to infer the latent variables. Li and Yu (2014) introduce Iterative Weighted Majority Voting (IWMV) to aggregate noisy crowd labels by iteratively estimating worker reliability, and show it approaches the oracle Maximum A Posteriori (MAP) solution.
@@
Rangi and Franceschetti (2018) utilize the bandits-with-knapsacks framework to dynamically estimate worker accuracy and allocate tasks in real time to maximize overall labeling quality within a budget. Another influential model is by Raykar et al. (2010), which jointly infers true labels and annotator reliabilities by modeling each worker’s confusion matrix. Through an expectation–maximization procedure, the method down-weights inconsistent annotators and yields more accurate aggregated labels without prior knowledge of worker quality. Our method parallels this online estimation but differs in that it leverages contextual embeddings and targets LLM ensembles rather than human annotators.
@@
Yang et al. (2023) Weighted voting Yes No Yes
Nguyen et al. (2024) Single-model routing Yes No Yes
Ding et al. (2024) Single-model routing Yes No Yes
Chen et al. (2024) Re-querying No No No
Li et al. (2024a) Re-querying No No Yes
Li and Yu (2014) Crowd aggregation No Yes No
Rangi and Franceschetti (2018) Crowd assignment No Yes No
Raykar et al. (2010) Crowd aggregation No Yes No
@@ 2 Problem Statement
To serve as a baseline, we introduce the following weighted majority voting scheme. In this scheme, the predicted label $y_{t}$ for instance $x_{t}$ is determined by aggregating the votes of all $K$ LLMs using:
$y_{t}=\arg\max_{m\in[M]}\sum_{i=1}^{K}\omega_{\text{def},i}(t)\cdot{\mathds{1}\left\{y_{i,t}=m\right\}},$ (1)
where ${\mathds{1}\left\{\cdot\right\}}$ is the indicator function that returns 1 if the condition is true and 0 otherwise, and $y_{i,t}$ denotes the label outputted by LLM $l_{i}$ for instance $x_{t}$. Since the true label is not available in our setting, we use the empirical accuracy of model $l_{i}$ as the voting weight. Hence, $\omega_{\text{def},i}(t)=(\sum_{s=1}^{t-1}\mathbb{I}(y_{i,s}=y_{s}))/N_{i,t},$ where $N_{i,t}$ denotes the number of times LLM $l_{i}$ has been queried up to round $t$, and $y_{s}$ is the predicted label for round $s$. The pseudocode for this scheme, which we refer to as the baseline method, is provided in Algorithm 2 in Appendix C.
@@
We will leverage the following result to estimate the confidence of the majority vote label based on the $L_{i,t}$ and $\omega_{i}(t)$ values when majority voting is performed over a subset $\mathcal{A}$ of LLMs.
@@
Finally, we introduce a user-specified parameter $k_{\min}$, which enforces a floor on the number of LLMs queried per instance. By requiring at least $k_{\min}$ votes in every round, this constraint further safeguards label quality—ensuring that no annotation is based on fewer than $k_{\min}$ model predictions.
@@ 3 The CaMVo Algorithm
We propose a novel algorithm, Cost-aware Majority Voting (CaMVo), for efficient dataset labeling with large language models (LLMs). CaMVo aims to select a cost-effective subset of LLMs for each input instance by leveraging contextual embeddings to estimate a lower confidence bound on the probability that each model will produce a correct label. These bounds are computed using a LinUCB-based framework (Li et al., 2010). Based on this information, CaMVo identifies a subset of LLMs such that the confidence of their weighted majority vote exceeds a user-specified threshold $\delta$. If no such subset exists, the algorithm defaults to querying all available models. The pseudo-code of CaMVo is provided in Algorithm 1, and consists of six main steps described below.
@@
1:Input: Set of LLMs $[K]$, cost per token $\rho_{i}$ for each LLM $i$, embedding model $\text{Emb}(\cdot)$, confidence threshold $\delta$, LinUCB regularization parameter $\lambda_{L}$, exploration parameter $\alpha$, regularization parameter $\lambda_{R}$
2:Initialize: $A_{i}\leftarrow\lambda_{L}I_{d},~{}\bm{b}_{i}\leftarrow 0_{d},~{}\forall i\in[K]$
@@
9:         $L_{i,t}\leftarrow\frac{\bar{L}_{i,t}\cdot N_{i,t}+\lambda_{R}\cdot\log(t+1)/2}{N_{i,t}+\lambda_{R}\cdot\log(t+1)}$
@@
14:     $y_{t}\leftarrow\arg\max_{m}\sum_{i\in\mathcal{A}_{t},y_{i,t}=m}\omega_{i}(t)$
@@
17:              $r_{i,t}\leftarrow{\mathds{1}\left\{y_{i,t}=y_{t}\right\}}$
18:              Update: $A_{i,t}\leftarrow A_{i,t-1}+\mathbf{e}_{t}\mathbf{e}_{t}^{\top}$
19:              Update: $\bm{b}_{i,t}\leftarrow\bm{b}_{i,t-1}+r_{i,t}\mathbf{e}_{t}$
20:              Update: $\mu_{i,t}\leftarrow\frac{\sum_{s=1}^{t}{\mathds{1}\left\{y_{i,s}=y_{s}\right\}}}{N_{i,t}}$
21:              Update the parameters of $\text{Est}_{i}$
22:         end for
23:     end if
24:end for
@@
For each LLM $l_{i}$, CaMVo maintains a matrix $A_{i}\in\mathbb{R}^{d\times d}$ and vector $\bm{b}_{i}\in\mathbb{R}^{d}$, initialized as $A_{i}=\lambda_{L}I_{d}$, $\bm{b}_{i}=\bm{0}$, where $\lambda_{L}>0$ is a user-defined parameter. Given $\bm{e}_{t}=\text{Emb}(x_{t})$, the estimated confidence, and its confidence bound is computed as:
@@
Given the inherent probabilistic nature of LLM outputs, the estimated confidence score may not reliably indicate the correctness of a label. To address this, we introduce a Bayesian estimator $\text{Est}_{i}(\cdot)$ that models the posterior probability that $l_{i}$’s prediction is correct, conditioned on this confidence. First, we define a latent variable $h_{i,t}=\mathds{1}\{y_{i,t}=y_{t}\}$, where $y_{t}$ is the assigned label. Note that ideally, the true label should be used instead of $y_{t}$, but since we do not have access to the true label, we instead use $y_{t}$. We model the conditional likelihood of $q_{i,t}(\bm{e}_{t})$, given the latent variable $h_{i,t}$, as a Beta-distributed random variable:
@@
Further, to model $\mathbb{P}(h_{i,t}=1)$, we use the empirical historical accuracy of $l_{i}$, $\mu_{i,t-1}$, which captures the accuracy of LLM $l_{i}$ relative to the predicted labels up to round $t-1$. Similarly, $\mathbb{P}(h_{i,t}=0)=1-\mu_{i,t-1}$. Applying Bayes’ rule with these models, the posterior probability is modeled as:
@@
In the absence of ground-truth labels, empirical accuracy estimates in majority voting can overfit, resulting in overconfident weights and biased aggregation. To address this issue, we regularize the LCB of the estimated confidence of LLM $l_{i}$ using Laplace smoothing:
@@
where $\lambda_{R}>0$ is a user-defined regularization parameter that controls the strength of smoothing.
@@
where $c_{\mathcal{A}}(t)=\sum_{i\in\mathcal{A}}\rho_{i}\cdot H_{i}(x_{t}),$ and $H_{i}(x_{t})$ is the token count of the query to the LLM $l_{i}$ for data instance $x_{t}$ under the tokenization method of LLM $l_{i}$. Since we expect only one token as output (which will be the label for the data instance $x_{t}$), we ignore the output tokens. Note that as in Lemma 2.1, we assume that the outputs of LLMs are conditionally independent given the data instance. Also note that if no such $\mathcal{A}$ exists, CaMVo defaults to querying all LLMs.
@@
We query the LLMs in $\mathcal{A}_{t}$ and receive their responses. The label for $x_{t}$ can be assigned via weighted majority vote using $y_{t}=\arg\max_{m\in[M]}\sum_{i\in\mathcal{A}_{t}}\omega_{i}(t)\cdot{\mathds{1}\left\{y_{i,t}=m\right\}}$.
@@
If $|A_{i,t}|>1$, for each $l_{i}\in\mathcal{A}_{t}$, CaMVo updates $A_{i,t}$, $\bm{b}_{i,t}$, and $\mu_{i,t}$, which is the empirical mean accuracy, as:
$A_{i,t}\leftarrow A_{i,t-1}+\bm{e}_{t}\bm{e}_{t}^{\top},\quad\bm{b}_{i,t}\leftarrow\bm{b}_{i,t-1}+r_{i,t}\bm{e}_{t},\quad\mu_{i,t}\leftarrow\frac{\sum_{s=1}^{t}\mathds{1}\{y_{i,s}=y_{s}\}}{N_{i,t}}$
where $r_{i,t}=\mathds{1}\{y_{i,t}=y_{t}\}$. Note that parameters are not updated when $|A_{i,t}|=1$ as the reward will always be $1$ in that case. The Beta distribution parameters $(\alpha_{i,h},\beta_{i,h})$ can be updated via maximum likelihood estimation or a method-of-moments approximation based on the mean and variance of past confidence scores. These approaches are discussed in Appendix B.

### v2

@@ Abstract
footnotetext: 39th Conference on Neural Information Processing Systems (NeurIPS 2025).
@@ 1 Introduction
The rapid proliferation of data across domains has created an urgent need for accurate, large‐scale annotation pipelines. While human experts and crowd workers have been the gold standard for dataset labeling, manual annotation is notoriously slow, expensive, and prone to inter‐annotator inconsistency petrovic2020crowdsourcing. As machine learning models become increasingly sophisticated, their demand for high‐quality, richly labeled datasets only intensifies, exacerbating this bottleneck.
Recent advances in large language models (LLMs) offer a promising remedy: by leveraging transformer‐based architectures such as GPT, it is now possible to automate much of the labeling workload. LLMs excel at natural‐language understanding, reasoning, and contextual inference, enabling rapid generation of annotations with minimal human effort naveed2023comprehensive.
However, relying on a single LLM introduces issues of biases inherited from its training data and stochastic variability across repeated queries, undermining reliability and reproducibility errica2024did; li2024showing. A common strategy to bolster label quality is ensembling: querying multiple LLMs, or multiple samples from the same model, and aggregating their outputs via majority voting. This reduces hallucinations and offsets the bias of individual models but substantially increases cost, as each additional model increases latency and compute expenditure yang2023one. In practice, querying every available LLM for each instance is often wasteful and unnecessary.
In this paper, we address this trade‐off by adaptively selecting a subset of LLMs for majority voting on each input, achieving comparable accuracy to full‐ensemble voting while dramatically cutting cost. Unlike prior work on LLM weight optimization or query routing—which presumes access to ground‐truth labels or a pre‐trained routing model chen2024more; nguyen2024metallm; ding2024hybrid—our method operates online, without any held‐out training set or ground truth.
@@
    Cost‐aware Majority Voting (CaMVo). We propose CaMVo, an algorithm that combines a LinUCB‐style contextual bandit with a Bayesian Beta‐mixture confidence estimator. For each candidate LLM, CaMVo computes a lower confidence bound on its correctness probability given the input’s embedding, then selects the smallest‐cost subset whose aggregated confidence exceeds a user‐specified threshold $\delta$.
@@ 1.1 Related Work
Aggregating outputs from multiple LLMs (or repeated queries to a single LLM) via majority voting has become a popular strategy to boost annotation reliability. chen2024more analyze the effect of repeated queries to a single model and observe a non‐monotonic accuracy curve: performance improves initially but degrades beyond an optimal number of calls due to task heterogeneity. While additional LM calls enhance accuracy on easier queries, they may introduce noise or inconsistency that degrades performance on more challenging ones. To address this, the authors propose a scaling model that predicts the optimal number of LM calls required to maximize aggregate performance. trad2024ensemble compare re‐querying and multi‐model ensembles, showing that ensemble strategies are most effective when individual models or prompts exhibit comparable performance levels, and ensemble gains may diminish when individual model accuracies diverge.
yang2023one propose a weighted majority‐voting ensemble for medical QA, combining dynamic weight adjustment with clustering‐based model selection. However, their approach relies on offline training data and focuses solely on improving accuracy, without accounting for the cost of querying. In contrast, our approach selects a cost‐effective subset of heterogeneous LLMs online, without any pre‐training.
@@
Query routing addresses the problem of selecting a single LLM per query to optimize cost or latency under performance constraints. nguyen2024metallm cast LLM selection as a contextual bandit problem, training offline on labeled data to learn a routing policy that maps query embeddings to a single optimal LLM under a total budget constraint of $b$. ding2024hybrid train a router to distinguish “easy” versus “hard” queries, sending easy tasks to local LLMs and hard ones to cloud APIs. Unlike these methods, our method operates in a fully online setting without access to labeled training data, updating the model dynamically during the labeling process. Moreover, rather than routing to a single LLM, we select a cost-efficient subset for majority voting.
@@
Estimating model confidence can guide automatic label selection. kadavath2022language explore how a language model’s own uncertainty estimates can serve as predictors of answer correctness, and find that the cumulative log‐probability the model assigns to its generated token sequence correlates strongly with factual accuracy across diverse benchmarks. li2024showing generate code multiple times and use output similarity as a proxy for confidence, choosing the most consistent result. While these works focus on self‐consistency of a single model, we estimate a probabilistic lower bound on each LLM’s correctness via a Bayesian Beta‐mixture model, incorporating both past performance and context.
@@
Beyond simple voting, weighted schemes assign each annotator or model a reliability score such as the accuracy of the annotator, or the label confidence reported by the annotator. One notable approach is the GLAD model by whitehill2009whose, which formulates weighted majority voting as a probabilistic inference problem over annotator expertise and task difficulty. GLAD jointly estimates per‐annotator reliability and per‐item ambiguity via a generative model, using an EM algorithm to infer the latent variables. li2014error introduce Iterative Weighted Majority Voting (IWMV) to aggregate noisy crowd labels by iteratively estimating worker reliability, and show it approaches the oracle Maximum A Posteriori (MAP) solution.
@@
rangi2018multi utilize the bandits‐with‐knapsacks framework to dynamically estimate worker accuracy and allocate tasks in real time to maximize overall labeling quality within a budget. Another influential model is by raykar2010learning, which jointly infers true labels and annotator reliabilities by modeling each worker’s confusion matrix. Through an expectation–maximization procedure, the method down‐weights inconsistent annotators and yields more accurate aggregated labels without prior knowledge of worker quality. Our method parallels this online estimation but differs in that it leverages contextual embeddings and targets LLM ensembles rather than human annotators.
@@
yang2023one Weighted voting Yes No Yes
nguyen2024metallm Single‐model routing Yes No Yes
ding2024hybrid Single‐model routing Yes No Yes
chen2024more Re‐querying No No No
li2024showing Re‐querying No No Yes
li2014error Crowd aggregation No Yes No
rangi2018multi Crowd assignment No Yes No
raykar2010learning Crowd aggregation No Yes No
@@ 2 Problem Statement
To serve as a baseline, we introduce the following weighted majority voting scheme. In this scheme, the predicted label $\hat{y}_{t}$ for instance $x_{t}$ is determined by aggregating the votes of all $K$ LLMs using:
$\hat{y}_{t}=\arg\max_{m\in[M]}\sum_{i=1}^{K}\omega_{\text{def},i}(t)\cdot{\mathds{1}\left\{y_{i,t}=m\right\}},$ (1)
where $\mathds{1}\left\{\cdot\right\}$} is the indicator function that returns 1 if the condition is true and 0 otherwise, and $y_{i,t}$ denotes the label outputted by LLM $l_{i}$ for instance $x_{t}$. Since the true label is not available in our setting, we use the empirical accuracy of model $l_{i}$ relative to the predicted labels as the voting weight. Hence, $\omega_{\text{def},i}(t)=(\sum_{s=1}^{t-1}\mathbb{I}(y_{i,s}=\hat{y}_{s}))/N_{i,t},$ where $N_{i,t}$ denotes the number of times LLM $l_{i}$ has been queried up to round $t$, and $\hat{y}_{s}$ is the predicted label for round $s$. The pseudocode for this scheme, which we refer to as the baseline method, is provided in Algorithm 2 in Appendix C.
@@
We will leverage the following result to estimate the confidence of the majority vote label based on the $L_{i,t}$ and $\omega_{i,t}$ values when majority voting is performed over a subset $\mathcal{A}$ of LLMs.
@@
For practical purposes and computational efficiency, this expression may be approximated as $\delta_{\mathcal{A}}(\bm{L},\bm{\omega})\approx 1-F_{\text{Beta}}\left(0.5;\,W_{L,\mathcal{A}},\,W_{\mathcal{A}}-W_{L,\mathcal{A}}\right),$ where $F_{\text{Beta}}(x;\alpha,\beta)$ is the CDF of a $\text{Beta}(\alpha,\beta)$ distribution evaluated at point $x$, $W_{L,\mathcal{A}}=\sum_{i\in\mathcal{A}}\omega_{i}\cdot L_{i}$, and $W_{\mathcal{A}}=\sum_{i\in\mathcal{A}}\omega_{i}$.
Note that Lemma 2.1 can be extended to remove the conditional-independence assumption using Bahadur’s model, which characterizes the joint distribution of binary variables using their marginal probabilities, second-order dependence terms known as Bahadur parameters, and higher-order interaction terms (bahadur1961classification). This framework allows the computation of majority vote confidence by summing the probabilities of all outcomes where the majority label is correct. However, the absence of a closed-form expression and the exponential growth in computational complexity with the number of variables limit its practical use. In practice, Monte Carlo–based simulation methods can approximate the majority-vote confidence by generating correlated LLM outputs; we describe and analyze CaMVo with one such approach in §G.
Finally, we introduce a user‐specified parameter $k_{\min}$, which enforces a floor on the number of LLMs queried per instance. By requiring at least $k_{\min}$ votes in every round, this constraint further safeguards label quality—ensuring that no annotation is based on fewer than $k_{\min}$ model predictions.
@@ 3 The CaMVo Algorithm
We propose a novel algorithm, Cost-aware Majority Voting (CaMVo), for efficient dataset labeling with large language models (LLMs). CaMVo aims to select a cost-effective subset of LLMs for each input instance by leveraging contextual embeddings to estimate a lower confidence bound on the probability that each model will produce a correct label. These bounds are computed using a LinUCB-based framework (li2010contextual). Based on this information, CaMVo identifies a subset of LLMs such that the confidence of their weighted majority vote exceeds a user-specified threshold $\delta$. If no such subset exists, the algorithm defaults to querying all available models. The pseudo-code of CaMVo is provided in Algorithm 1, and consists of six main steps described below.
@@
1:Input: Set of LLMs $[K]$, cost per token $\rho_{i}$ for each LLM $i$, embedding model $\text{Emb}(\cdot)$, confidence threshold $\delta$, LinUCB regularization parameter $\lambda_{L}$, exploration parameter $\alpha$, regularization parameter $\lambda_{R}$, $k_{\min}$
[…]
@@
9:   $L_{i,t}\leftarrow\frac{\bar{L}_{i,t}\cdot N_{i,t-1}+\lambda_{R}\cdot\log(t+1)/2}{N_{i,t-1}+\lambda_{R}\cdot\log(t+1)}$
@@
14:  $\hat{y}_{t}\leftarrow\arg\max_{m}\sum_{i\in\mathcal{A}_{t},y_{i,t}=m}\omega_{i}(t)$
@@
17:     $r_{i,t}\leftarrow{\mathds{1}\left\{y_{i,t}=\hat{y}_{t}\right\}}$
18:     Update: $N_{i,t}\leftarrow N_{i,t-1}+1$
19:     Update: $A_{i,t}\leftarrow A_{i,t-1}+\mathbf{e}_{t}\mathbf{e}_{t}^{\top}$
20:     Update: $\bm{b}_{i,t}\leftarrow\bm{b}_{i,t-1}+r_{i,t}\mathbf{e}_{t}$
21:     Update: $\mu_{i,t}\leftarrow\frac{\sum_{s=1}^{t}{\mathds{1}\left\{y_{i,s}=\hat{y}_{s}\right\}}}{N_{i,t}}$
22:     Update the parameters of $\text{Est}_{i}$
23:   end for
24:  end if
25:end for
@@
For each LLM $l_{i}$, CaMVo maintains a matrix $A_{i}\in\mathbb{R}^{d\times d}$ and vector $\bm{b}_{i}\in\mathbb{R}^{d}$, initialized as $A_{i,0}=\lambda_{L}I_{d}$, $\bm{b}_{i,0}=\bm{0}$, where $\lambda_{L}>0$ is a user-defined parameter. Given $\bm{e}_{t}=\text{Emb}(x_{t})$, the estimated confidence, and its confidence bound is computed as:
@@
Since this is an LCB of a probability, we clip it between 0 and 1 when it is out of range [0,1].
@@
Given the inherent probabilistic nature of LLM outputs, the estimated confidence score may not reliably indicate the correctness of a label. To address this, we introduce a Bayesian estimator $\text{Est}_{i}(\cdot)$ that models the posterior probability that $l_{i}$’s prediction is correct, conditioned on this confidence. First, we define a latent variable $h_{i,t}=\mathds{1}\{y_{i,t}=\hat{y}_{t}\}$, where $\hat{y}_{t}$ is the assigned label. Note that ideally, the true label should be used instead of $\hat{y}_{t}$, but since we do not have access to the true label, we instead use $\hat{y}_{t}$. We model the conditional likelihood of $q_{i,t}(\bm{e}_{t})$, given the latent variable $h_{i,t}$, as a Beta-distributed random variable:
@@
Further, to model $\mathbb{P}(h_{i,t}=1)$, we use the empirical historical relative accuracy of $l_{i}$, $\mu_{i,t-1}$, which captures the accuracy of LLM $l_{i}$ relative to the predicted labels up to round t1t-1. Similarly, $\mathbb{P}(h_{i,t}=0)=1-\mu_{i,t-1}$. Applying Bayes’ rule with these models, the posterior probability is modeled as:
@@
In the absence of ground-truth labels, empirical relative accuracy estimates in majority voting can overfit, resulting in overconfident weights and biased aggregation. Further, since subset selection is based on these estimated confidences, early estimation errors can bias the selection process and lead to compounding errors over time. To address this issue, we regularize the LCB of the estimated confidence of LLM $l_{i}$ using Laplace smoothing:
@@
where $\lambda_{R}>0$ is a user-defined regularization parameter that controls the strength of smoothing. Laplace smoothing is chosen since it corresponds to using a uniform prior on the Beta estimator, and hence combines well with the Beta estimator. We use a $\log(t)$ term with the Laplace smoothing as compared to a constant term, as the $\log(t)$ term does not decay as quickly as the constant term and prevents overfitting in the long run.
@@
where $c_{\mathcal{A}}(t)=\sum_{i\in\mathcal{A}}\rho_{i}\cdot H_{i}(x_{t})$ , $\rho_{i}$ is the cost per token for LLM $i$, and $H_{i}(x_{t})$ is the token count of the query to the LLM $l_{i}$ for data instance $x_{t}$ under the tokenization method of LLM $l_{i}$. Since we expect only one token as output (which will be the label for the data instance $x_{t}$), we ignore the output tokens. Note that as in Lemma 2.1, we assume that the outputs of LLMs are conditionally independent given the data instance. We analyze the correlated version of CaMVo that does not have this assumption in Appendix G. Also note that if no such $\mathcal{A}$ exists, CaMVo defaults to querying all LLMs.
@@
We query the LLMs in $\mathcal{A}_{t}$ and receive their responses. The label for $x_{t}$ can be assigned via weighted majority vote using $\hat{y}_{t}=\arg\max_{m\in[M]}\sum_{i\in\mathcal{A}_{t}}\omega_{i}(t)\cdot{\mathds{1}\left\{y_{i,t}=m\right\}}$.
@@
If $|\mathcal{A}_{t}|>1$, for each $l_{i}\in\mathcal{A}_{t}$, CaMVo updates $A_{i,t}$, $\bm{b}_{i,t}$, and $\mu_{i,t}$, which is the empirical relative mean accuracy, as:
$A_{i,t}\leftarrow A_{i,t-1}+\bm{e}_{t}\bm{e}_{t}^{\top},\quad\bm{b}_{i,t}\leftarrow\bm{b}_{i,t-1}+r_{i,t}\bm{e}_{t},\quad\mu_{i,t}\leftarrow\frac{\sum_{s=1}^{t}\mathds{1}\{y_{i,s}=\hat{y}_{s}\}}{N_{i,t}}$
where $r_{i,t}=\mathds{1}\{y_{i,t}=\hat{y}_{t}\}$. Note that parameters are not updated when $|\mathcal{A}_{t}|=1$ as the reward will always be $1$ in that case. The Beta distribution parameters $(\alpha_{i,h},\beta_{i,h})$ can be updated via maximum likelihood estimation or a method-of-moments approximation based on the mean and variance of past confidence scores. These approaches are discussed in Appendix B.